40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8347     Entry time: Tue Mar 26 00:06:37 2013
Author: Jenne 
Type: Update 
Category: Locking 
Subject: AS beam put back on PD 

[Jenne, Gabriele]

We aligned MICH (first locked Yarm, but didn't optimize since we don't have TRY, then locked Xarm, then aligned MICH), but there was no beam on AS55.  We went out to check, and the beam was almost not hitting the small steering mirror between AS55.  We adjusted the BS splitting the beam between camera and PD, and got the beam back on AS55. We could then lock MICH.

We also futzed with the REFL55 phase to get PRCL stuff in I, and MICH stuff in Q.  The procedure was to align PRMI, then kick PRM in pos, and adjust the phase so we got signal mostly in I after the kick.  We started at the original value of 60deg, but are leaving it at -20deg.

ELOG V3.1.3-